# UQE: A Query Engine for Unstructured Databases

### v1

@@ 6.1.3 Abstraction and aggregation
This task abstracts the intrinsics of each row, and then performs semantics-based GROUP BY, grouping the common instrinsics across all rows. Finally, it provides aggregated statistics over each group: {minted}[fontsize=]mysql SELECT derived_attribute, COUNT(*) FROM table GROUP BY "extract an abstract intrinsic attribute specified in natural language" AS derived_attribute LIMIT 10 The challenging problems in this task are (i) building a taxonomy with good coverage, and (2) bias and variance reduction for groups with small population. The result of this query is a list of tuples of derived attributes and their number of occurrences in the dataset. We use the earth mover’s distance (EMD (Pele and Werman, 2009)) as the evaluation metric to compare the extracted tuples and ground-truth tuples. The distance between a pair of attributes is defined by one minus the cosine similarity of their text embeddings. We can see from Table 3 that UQE consistently outperforms baselines while achieving much lower cost. We also show in Appendix C with more qualitative results comparisons.
6.2 Ablation results
@@ 6.2.1 Variance of different sampling approaches for aggregation queries
To decouple the variance introduced by the algorithm and the bias introduced by the LLM based predictors, here we use the ground-truth label as the predictive result and focus on the effectiveness of variance reduction. Figure 4 shows the box plot of different sampling methods. We can see using stratified sampling over the embeddings of unstructured content achieves significant lower variance compared to the uniform random sampling. Also both of these achieve similar expected values, which also justifies the correctness or unbiasness.
@@ 7 Conclusion
This paper proposed an unstructured query engine that leverages 1) the flexibility of LLMs for data understanding; 2) the advances in sampling and online learning for efficient data scanning; 3) and the compiler that bridges these algorithmic workflows with LLMs. We demonstrated its efficiency and accuracy over three analytic tasks on four datasets with two different modailities. However the current work is still very limited in terms of 1) the semantics it lacks, including table join and other types of aggregations; 2) an automated selection of LLMs and sampling configurations; 3) and scaling to even larger databases. We hope to investigate these further in future works.
@@ C.1 Datasets
  • User review mining. We use IMDB (Maas et al., 2011) for semantic analysis of user sentiment. The entire dataset contains 50K highly polar movie reviews with positive and negative sentiment labels.
  • Goal-oriented (customer service) dialogue systems. We use two datasets for this category, including 1) Action-Based Conversations Dataset (ABCD (Chen et al., 2021)) for intent identification, which has 10,042 dialogs with 10 distinct user intents requiring unique sequences of actions constrained by policies to achieve task success; and 2) AirDialog (Wei et al., 2018) for conversation outcome understanding. It contains 402,037 goal-oriented conversation on flight booking with 5 possible ground-truth states.
  • Image understanding. We use Clevr (Johnson et al., 2017) dataset for multimodal data understanding and retrieving. Specifically we use the val split of the dataset, which contains 15,000 images of objects containing different number of cylinders, cubes and spheres with different sizes/colors. We downsample the image to size no more than $128\times 128$, and only feed the images to the LLMs while holding out scene metadata for evaluation only.
@@
Dataset Conditions percentage of occurance in the data
@@ C.2 Parameter and experiment setup
C.3 Latency
@@
C.4 Group By qualitative results
For single_item_query in ABCD, the items menteiond in the dialogs found by UQE is:
@@ C.4 Group By qualitative results
For account_access in ABCD, the issues menteiond in the dialogs found by UQE is:
@@ D.1.1 Task: Conditional aggregation
Read the following movie reviews, and categorize them into either positive or negative class, depending on the sentiment of the review. If the movie has a mixed sentiment, try your best to classify into positive or negative class based on the overall sentiment.In the end, please just output a single number, which is [the total number of positive reviews].
@@
[Review 0]: Some TV programs continue into embarrassment (my beloved ’X-Files’ comes to mind.)…
[Review 1]: The tale of the titular Adam (Mark O’ Halloran) and Paul (Tom Murphy), …
@@
The following dialogs between a customer service agent and a customer. Dialogs start by headers such as **Dialog 1**, **Dialog 2**, and so on. Your task is to classify whether the dialog content is about the theme "account access issue". Then count how many dialogs are talking about this theme. In the end, output the count as a single number. Here is more detailed explanation about Theme "<THEME>". Be sure to use this information when you classify. Theme "<THEME>" dialogs content is about THEME_EXPLANATION. Please perform thorough analysis for each of the dialog. In the end, please **only** output a single number, which is [the total number of dialogs] that talks about the theme "<THEME>".
In the paper we use the following themes and corresponding explanations:
@@
  • <THEME> = requesting detailed specifications of a certain item sold on the website.
  • <THEME_EXPLANATION> = the customer needed help with the detailed specifications of a certain item sold on the website, such as inquiries about detailed information of a specific retail item about materials, whether it shrinks, stock availability, etc., but NOT inquiries about promotions, order status, shipping status, questions about how to use the website to purchase an item, or difficulties on using the website to add to cart or purchase an item, or inquiries about subscription.
@@
[customer]: Hello my name is Alessandro Phoenix and I need to make sure the shipping cost is included on my order …
@@
[change]: the agent changed the existing flight reservation of the customer and successfully found a new one;
[no_reservation]: the customer wants to change or cancel the flight but1037 there is no valid reservation under this customer;
@@
customer: Can you please find a flight from DFW to SEA? …
@@
customer: I want to celebrate Thanks giving day 11/23 with my friends at New York. Can you book a ticket for me? …
@@
image_0: <base64_encoded_image>
image_1: <base64_encoded_image> …
@@ D.1.2 Task: Semantic retrieval
Read the following movie reviews, and list the indices of reviews with positive sentiment. If the movie has a mixed sentiment, try your best to classify into positive or negative class based on the overall sentiment.In the end, please only output a list of indices in the format of [review_3, review_7, …]
@@
[Review 0]: Some TV programs continue into embarrassment (my beloved ’X-Files’ comes to mind.)…
[Review 1]: The tale of the titular Adam (Mark O’ Halloran) and Paul (Tom Murphy), …
@@
The following are dialogs between a customer service agent and a customer. Dialogs start by headers such as **dialog_1**, **dialog_2**, and so on. Your task is to find out the dialogs where <CONDITION>. Please only output a list of indices in the format of [dialog_3, dialog_7, …]
where the <CONDITION> can be one of the following:
@@
[customer]: Hi. I am very frustrated because I am trying to use your website and it is running SO slowly! …
@@
[customer]: i wanted to know if you’d be able to tell me the arm length on a shirt i’m thinking of buying? …
@@
System prompt
The following are dialogs between a airline ticketing agent and a customer. Dialogs start with headers such as **dialog_1**, **dialog_2**, and so on. Your task is to find out the dialogs where <CONDITION>. Please only output a list of indices in the format of [dialog_3, dialog_7, …]
where the <CONDITION> can be one of the following:
@@
customer: Can you please find a flight from DFW to SEA? …
**dialog_1**:
@@
customer: I want to celebrate Thanks giving day 11/23 with my friends at New York. Can you book a ticket for me? …
@@
Please read and parse the following images. Images start with labels such as **image_0**, **image_1**, and so on. Your task is to find out the images where <CONDITION>. Please only output a list of indices in the format of [image_3, image_7, …]
where the <CONDITION> can be one of the following:
@@
image_1: <base64_encoded_image> … Given above, the relevant images are:
@@ D.1.3 Task: Abstraction and aggregation
The following are dialogs between a customer service agent and a customer. Dialogs start with headers such as **dialog_1**, **dialog_2**, and so on. Your task is to analyze all the dialogs, and summarize "<ABSTRACT_ATTRIBUTE>" into groups. Please output the table of your analysis, in the format of pairs of ("<ABSTRACT_ATTRIBUTE>", number_of_dialogs belong to that). Specifically in the format as:
@@
where the <ABSTRACT_ATTRIBUTE> can be one of the following:
  • <ABSTRACT_ATTRIBUTE> = the type of account access issue
  • <ABSTRACT_ATTRIBUTE> = the single item involved in the dialog
@@
[customer]: hey think i mixed up or forgot which username I’m in with you guys as …
@@
[agent]: Hi there, thanks for contacting Acme! How can I help you? …
@@
The following are dialogs between a airline ticketing agent and a customer. Dialogs start with headers such as **dialog_1**, **dialog_2**, and so on. Your task is to analyze all the dialogs, and summarize "<ABSTRACT_ATTRIBUTE>" into groups. Please output the table of your analysis, in the format of pairs of ("<ABSTRACT_ATTRIBUTE>", number_of_dialogs belong to that). Specifically in the format as:
@@
group 2,number_of_dialogs …
where <ABSTRACT_ATTRIBUTE> = the outcome of the dialog
@@
customer: I need to book a flight ticket from DEN to EWR to enjoy music festivals. …
@@
agent: Hello, how may I help you? …
@@ D.2.1 Task: Conditional aggregation, Semantic retrieval
[Movie review]: May I please have my \$13.00 back? I would have rather watched "Hydro- Electric Power Comes to North America"…
@@
[action]: Searching the FAQ pages …
@@ D.2.2 Task: Abstraction and aggregation
    The following are dialogs between a customer service agent and a customer. Dialogs start with headers such as **dialog_1**, **dialog_2**, and so on. Your task is to analyze all the dialogs, and summarize "<ABSTRACT_ATTRIBUTE>" into groups. Please output the table of your analysis, in the format of pairs of ("<ABSTRACT_ATTRIBUTE>", number_of_dialogs belong to that). Specifically in the format as:
@@
    group 2,number_of_dialogs …
    with
    • <ABSTRACT_ATTRIBUTE> = the type of account access issue
    • <ABSTRACT_ATTRIBUTE> = the single item involved in the dialog
@@
    [customer]: hey think i mixed up or forgot which username I’m in with you guys as …
@@
    [agent]: Hi there, thanks for contacting Acme! How can I help you? …
@@
    Read the given airline ticketing dialog between an agent and a customer, and classify issue, into one or several categories below. Here are the description of the 3 categories:
@@
    [1]: Forgot PasswordOnly
    [2]: Two-factor authentication issues
    reply the index of the category, separated by ",". Here is the example format: [0, 3]
@@
    The following are dialogs between a airline ticketing agent and a customer. Dialogs start with headers such as **dialog_1**, **dialog_2**, and so on. Your task is to analyze all the dialogs, and summarize "<ABSTRACT_ATTRIBUTE>" into groups. Please output the table of your analysis, in the format of pairs of ("<ABSTRACT_ATTRIBUTE>", number_of_dialogs belong to that). Specifically in the format as:
@@
    group 2,number_of_dialogs …
    with <ABSTRACT_ATTRIBUTE> = the outcome of the dialog
@@
    customer: I need to book a flight ticket from DEN to EWR to enjoy music festivals. …
@@
    agent: Hello, how may I help you? …
@@
    Read the given airline ticketing dialog between the customer and the agent, and classify outcome, into one or several categories below. Here are the description of the 2 categories: [0]: Reservation cancelled [1]: Ticket booked
    Only reply the index of the category, separated by ",". Here is the example format: [0, 3]
@@
    Here is the airline ticketing dialog:
    customer: Hello

### v2

@@ 6.1.3 Abstraction and aggregation
This task abstracts the intrinsics of each row, and then performs semantics-based GROUP BY, grouping the common intrinsics across all rows. Finally, it provides aggregated statistics over each group: {minted}[fontsize=]mysql SELECT derived_attribute, COUNT(*) FROM table GROUP BY "extract an abstract intrinsic attribute specified in natural language" AS derived_attribute LIMIT 10 The challenging problems in this task are (i) building a taxonomy with good coverage, and (2) bias and variance reduction for groups with small population. The result of this query is a list of tuples of derived attributes and their number of occurrences in the dataset. We use the earth mover’s distance (EMD [30]) as the evaluation metric to compare the extracted tuples and ground-truth tuples. The distance between a pair of attributes is defined by one minus the cosine similarity of their text embeddings. We can see from Table 3 that UQE consistently outperforms baselines while achieving much lower cost. We also show in Appendix C with more qualitative results comparisons.
6.2 Ablation studies
We study the effectiveness of UQE for aggregation queries in Section 6.2.1 and non-aggregation queries in Section 6.2.2, and the quality/cost trade-off of UQE in Section 6.2.3. In appendix we provide more results on other modalities C.3, consistency C.4 and latency C.5.
@@ 6.2.1 Variance of different sampling approaches for aggregation queries
To decouple the variance introduced by the algorithm and the bias introduced by the LLM based predictors, here we use the ground-truth label as the predictive result and focus on the effectiveness of variance reduction. Figure 4 shows the box plot of different sampling methods. We can see using stratified sampling over the embeddings of unstructured content achieves significant lower variance compared to the uniform random sampling. Also both of these achieve similar expected values, which also justifies the correctness or unbiasedness.
@@ 6.2.2 Efficiency of online learning for non-aggregation queries
6.2.3 Trade-off between cost/latency and accuracy
Table 4: Quality at different compute budget $B$.
Budget $B$ 256 128 64 32
Retrieval latency(s) 38.08 21.61 11.11 5.84
F1 score 0.978 0.974 0.921 0.828
Aggregation latency(s) - 5.83 4.28 2.93
Error - 5.75% 6.84% 8.29%
Table 5: Error of the aggregation operation on IMDB dataset under different budget $B$
Budget $B$ 512 256 128 64 32
UDF 8.11% 8.35% 13.67% - -
UQE - - 5.75% 6.84% 8.29%
Generally the larger compute budget $B$ the better quality UQE will get, and we verify this in Table 5 5. UQE can achieve pretty good quality even with very low budget, and notably compared to the baseline, it achieves similar quality with 16x reduction of the compute needed. We show more results in Section C.5 regarding the compute efficiency.
@@ 7 Conclusion
This paper proposed an unstructured query engine that leverages 1) the flexibility of LLMs for data understanding; 2) the advances in sampling and online learning for efficient data scanning; 3) and the compiler that bridges these algorithmic workflows with LLMs. We demonstrated its efficiency and accuracy over three analytic tasks on four datasets with two different modalities. However the current work is still very limited in terms of 1) the semantics it lacks, including table join and other types of aggregations; 2) an automated selection of LLMs and sampling configurations; 3) and scaling to even larger databases. We hope to investigate these further in future works.
Acknowledgments and Disclosure of Funding
We thank Carsten Binnig, Howie Xu, Ras Bodik, Xinyi Chen, and the anonymous reviewers for providing helpful discussion and suggestions.
@@ C.1 Datasets
  • User review mining. We use IMDB [27] for semantic analysis of user sentiment. The entire dataset contains 50K highly polar movie reviews with positive and negative sentiment labels.
  • Goal-oriented (customer service) dialogue systems. We use two datasets for this category, including 1) Action-Based Conversations Dataset (ABCD [8]) for intent identification, which has 10,042 dialogs with 10 distinct user intents requiring unique sequences of actions constrained by policies to achieve task success; and 2) AirDialog [41] for conversation outcome understanding. It contains 402,037 goal-oriented conversation on flight booking with 5 possible ground-truth states.
  • Image understanding. We use Clevr [23] dataset for multimodal data understanding and retrieving. Specifically we use the val split of the dataset, which contains 15,000 images of objects containing different number of cylinders, cubes and spheres with different sizes/colors. We down-sample the image to size no more than $128\times 128$, and only feed the images to the LLMs while holding out scene metadata for evaluation only.
@@
Dataset Conditions percentage of occurrence in the data
@@ C.2 Parameter and experiment setup
C.3 Audio modality
We use the audio MNIST [5] data for this experiment. This dataset contains 30k wav files from 60 speakers pronouncing digits 0-9. We perform the audio semantic retrieval experiments. The query is first converted to audio space using TTS and the corresponding audio embedding is used for MIPS search. We use Gemini Pro 1.5 as the backend model, as it supports the audio inputs nicely. As is shown in Table 7, the proposed UQE consistently does better than alternatives, and it also allows complex queries that require reasoning, while embedding based MIPS is limited to certain types of queries.
Table 7: Retrieval F1 score on AudioMnist dataset.
Query UDF-gemini MIPS UQE-gemini
"Three" 0.109 $\pm$ 0.022 0.445 0.839 $\pm$ 0.135
"A number whose square equals to itself" 0.259 $\pm$ 0.068 0.039 0.922 $\pm$ 0.024
"The minimum prime number" 0.107 $\pm$ 0.044 0 0.917 $\pm$ 0.025
C.4 Consistency of the execution results
While UQE is able to leverage the foundation models to analyze the unstructured databases directly, the execution result is not deterministic due to the nature of the stochasticity of the algorithm and the LLMs themselves. In our paper we aim at reducing the variance so as to improve the consistency. However when the backend LLM gets updated, it might also cause the potential inconsistency. Below we analyze the effect of different LLM backends.
Table 8: Performance of UQE with different LLM backends.
Task UDF UQE-haiku (in the main paper) UQE-gpt4o-mini-0718
Retrieval 0.505 $\pm$ 0.030 0.978 $\pm$ 0.003 0.956 $\pm$ 0.010
Aggregation 13.67% $\pm$ 6.24% 5.75% $\pm$ 3.43% 6.33% $\pm$ 4.71%
Table 8 shows the effect on the IMDB dataset, where we see little variation. Actually the difference caused by the model switching is much lower than using a worse query engine. Of course, this behavior shift would be task related, but with the advances of LLMs we believe this variation would converge and be more stable to different prompts in the future.
C.5 Latency
@@
C.6 Group By qualitative results
For single_item_query in ABCD, the items mentioned in the dialogs found by UQE is:
@@ C.6 Group By qualitative results
For account_access in ABCD, the issues mentioned in the dialogs found by UQE is:
@@ D.1.1 Task: Conditional aggregation
Read the following movie reviews, and categorize them into either positive or negative class, depending on the sentiment of the review.If the movie has a mixed sentiment, try your best to classify into positive or negative class based on the overall sentiment.In the end, please just output a single number, which is [the total number of positive reviews]
@@
[Review 0]: Some TV programs continue into embarrassment (my beloved X-Files comes to mind.)...
[Review 1]: The tale of the titular Adam (Mark O Halloran) and Paul (Tom Murphy), ...
@@
The following dialogs between a customer service agent and a customer. Dialogs start by headers such as **Dialog 1**, **Dialog 2**, and so on.
Your task is to classify whether the dialog content is about the theme "account access issue". Then count how many dialogs are talking about this theme. In the end, output the count as a single number.
Here is more detailed explanation about Theme "<THEME>". Be sure to use this information when you classify.
Theme "<THEME>" dialogs content is about THEME_EXPLANATION.
Please perform thorough analysis for each of the dialog. In the end, please **only** output a single number, which is [the total number of dialogs] that talks about the theme "account access issue".
@@
<THEME> = requesting detailed specifications of a certain item sold on the website
@@
<THEME_EXPLANATION> = the customer needed help with the detailed specifications of a certain item sold on the website, such as inquiries about detailed information of a specific retail item about materials, whether it shrinks, stock availability, etc., but NOT inquiries about promotions, order status, shipping status, questions about how to use the website to purchase an item, or difficulties on using the website to add to cart or purchase an item, or inquiries about subscription
@@
[customer]: Hello my name is Alessandro Phoenix and I need to make sure the shipping cost is included on my order
@@
[no_reservation]: the customer wants to change or cancel the flight 1037 but there is no valid reservation under this customer;
@@
customer: Can you please find a flight from DFW to SEA?
@@
customer: I want to celebrate Thanks giving day 11/23 with my friends
at New York. Can you book a ticket for me?
@@
image\_0: <base64\_encoded\_image>
image\_1: <base64\_encoded\_image>
@@ D.1.2 Task: Semantic retrieval
Read the following movie reviews, and list the indices of reviews with positive sentiment. If the movie has a mixed sentiment, try your best to classify into positive or negative class based on the overall sentiment.In the end, please only output a list of indices in the format of [review_3, review_7, ...]
@@
[Review 0]: Some TV programs continue into embarrassment (my beloved X-Files comes to mind.)...
[Review 1]: The tale of the titular Adam (Mark O Halloran) and Paul (Tom Murphy), ...
@@
The following are dialogs between a customer service agent and a customer. Dialogs start by headers such as **dialog_1**, **dialog_2**, and so on.
Your task is to find out the dialogs where <CONDITION>. Please only output a list of indices in the format of [dialog_3, dialog_7, ...]
@@
[customer]: Hi. I am very frustrated because I am trying to use your website and it is running SO slowly!
@@
[customer]: i wanted to know if youd be able to tell me the arm length on a shirt im thinking of buying?
@@
System prompt The following are dialogs between a airline ticketing agent and a customer. Dialogs start with headers such as **dialog_1**, **dialog_2**, and so on. Your task is to find out the dialogs where CONDITION. Please only output a list of indices in the format of [dialog_3, dialog_7, ...]
@@
customer: Can you please find a flight from DFW to SEA?
**Dialog 1**:
@@
customer: I want to celebrate Thanks giving day 11/23 with my friends
at New York. Can you book a ticket for me?
@@
Please read and parse the following images. Images start with labels such as **image_0**, **image_1**, and so on.
Your task is to find out the images where <CONDITION>. Please only output a list of indices in the format of [image_3, image_7, ...]
@@
image_1: <base64_encoded_image>
Given above, the relevant images are:
@@ D.1.3 Task: Abstraction and aggregation
The following are dialogs between a customer service agent and a customer. Dialogs start with headers such as **dialog_1**, **dialog_2**, and so on.
Your task is to analyze all the dialogs, and summarize "<ABSTRACT_ATTRIBUTE>" into groups. Please output the table of your analysis, in the format of pairs of ("<ABSTRACT_ATTRIBUTE>", number_of_dialogs belong to that). Specifically in the format as:
@@
<ABSTRACT_ATTRIBUTE> = the type of account access issue <ABSTRACT_ATTRIBUTE> = the single item involved in the dialog
@@
[customer]: hey think i mixed up or forgot which username Im in with you guys as
@@
[agent]: Hi there, thanks for contacting Acme! How can I help you?
@@
The following are dialogs between a airline ticketing agent and a customer. Dialogs start with headers such as **dialog_1**, **dialog_2**, and so on.
Your task is to analyze all the dialogs, and summarize "<ABSTRACT_ATTRIBUTE>" into groups. Please output the table of your analysis, in the format of pairs of ("<ABSTRACT_ATTRIBUTE>", number_of_dialogs belong to that). Specifically in the format as:
@@
group 2,number_of_dialogs
<ABSTRACT_ATTRIBUTE> = the outcome of the dialog
@@
customer: I need to book a flight ticket from DEN to EWR to enjoy music festivals.
@@
agent: Hello, how may I help you?
@@ D.2.1 Task: Conditional aggregation, Semantic retrieval
[Movie review]: May I please have my \$13.00 back? I would have rather watched "Hydro- Electric Power Comes to North America"...
@@
[action]: Searching the FAQ pages ...
@@ D.2.2 Task: Abstraction and aggregation
    The following are dialogs between a customer service agent and a customer. Dialogs start with headers such as **dialog_1**, **dialog_2**, and so on.
    Your task is to analyze all the dialogs, and summarize "<ABSTRACT_ATTRIBUTE>" into groups. Please output the table of your analysis, in the format of pairs of ("<ABSTRACT_ATTRIBUTE>", number_of_dialogs belong to that). Specifically in the format as:
@@
    group 2,number_of_dialogs
    <ABSTRACT_ATTRIBUTE> = the type of account access issue <ABSTRACT_ATTRIBUTE> = the single item involved in the dialog
@@
    [customer]: hey think i mixed up or forgot which username Im in with you guys as
@@
    [agent]: Hi there, thanks for contacting Acme! How can I help you?
@@
    Read the given airline ticketing dialog between an agent and a customer, and classify issue, into one or several categories below. Here are the description of the 2 categories:
@@
    [1]: Forgot PasswordOnly reply the index of the category, separated by ",". Here is the example format:
    [0, 3]
@@
    The following are dialogs between a airline ticketing agent and a customer. Dialogs start with headers such as **dialog_1**, **dialog_2**, and so on.
    Your task is to analyze all the dialogs, and summarize "<ABSTRACT_ATTRIBUTE>" into groups. Please output the table of your analysis, in the format of pairs of ("<ABSTRACT_ATTRIBUTE>", number_of_dialogs belong to that). Specifically in the format as:
@@
    group 2,number_of_dialogs
    <ABSTRACT_ATTRIBUTE> = the outcome of the dialog
@@
    customer: I need to book a flight ticket from DEN to EWR to enjoy music festivals.
@@
    agent: Hello, how may I help you?
@@
    Read the given airline ticketing dialog between the customer and the agent, and classify outcome, into one or several categories below. Here are the description of the 2 categories:
    [0]: Reservation cancelled
    [1]: Ticket bookedOnly reply the index of the category, separated by ",". Here is the example format:
    [0, 3]
@@
    Here is the airline ticketing dialog: customer: Hello